# A Reinforcement Learning Method to Factual and Counterfactual Explanations for Session-based Recommendation

### v1

@@ 5.4. Ablation Study (RQ3)
We conduct experiments to evaluate the innovative designs in FCESR: (1) complex reward functions and (2) the impact of positive and negative samples.
Therefore, for the reward function, we consider the following three variants: FCESR (w/o) rank, which excludes the rank reward; FCESR (w/o) spars, which excludes both sparsity reward and rank reward; and FCESR (w/) fe, which utilizes only the factual reward. For our experiments, we utilized the Baby dataset (a subset of Amazon product data) and the MovieLens dataset, representing e-commerce and movie recommendation scenarios respectively. Experiment results on both datasets using SRGNN model (e.g., similar trends can be observed with the rest SR models) are presented in Table 5. The results demonstrate that each component significantly contributes to overall explanation performance, validating the robustness of the reward design.
Besides, the results reveal a significant trade-off between the completeness and conciseness of explanations. On Baby, FCESR (w/o) spars shows that removing the sparsity reward leads to an increase in the size of the explanation subsets found. Although FCESR (w/) fe has the best PS (sufficiency) at 68.24%, its explanation length (avg_len) is too long, indicating that it may not find the minimal subset. This demonstrates the role of the sparsity reward in balancing explanation quality and conciseness. The complete FCESR model, with its high $F_{ns}$ (61.31%) and reasonable avg_len (2.93), achieves a better balance between completeness and conciseness. The significance of the factual reward is evident from the performance of FCESR (w/) fe, which employs only the factual reward. On Baby, FCESR (w/) fe shows high PS (68.24%) but low PN (56.99%), indicating that relying solely on factual information may result in less comprehensive explanations. This observation is further supported by results on MovieLens, where FCESR (w/) fe achieves a more balanced performance but still falls short in terms of $F_{ns}$ (overall performance) compared to the complete FCESR. These findings underscore the necessity of incorporating multiple reward components to generate explanations that are both accurate and comprehensive. While simplified variants may perform better on certain metrics, the complete FCESR achieves the best balance between explanation quality, conciseness, and comprehensiveness.
@@ 5.5. Sensitivity of Hyper-parameters (RQ4)
Table 6. Runtime performance on different datasets.
Dataset Runtime (seconds)
Baby 115.6
Beauty 218.2
MovieLens 253.2
@@ 5.6. Time Complexity Analysis (RQ5)
The experiments were conducted on a computational infrastructure running Ubuntu 18.04, with 8 CPU cores and 64GB of memory. Table 6 presents the average runtime per episode of our RL framework across three datasets. The total training time varies depending on the stopping criteria (reward convergence, network convergence, or maximum episode limit). The runtime performance exhibits correlations with both the average session length and the total number of sessions in each dataset, as our framework needs to traverse all items within a session during each episode. For instance, the Baby dataset exhibits notably shorter runtime compared to the other two datasets, primarily attributed to its shorter session length. Overall, the computational complexity remains moderate, with session length being the dominant factor affecting runtime performance.

### v2

@@ 5.4. Ablation Study (RQ3)
We evaluate the key designs of FCESR, namely complex reward functions and the use of positive and negative samples. For the reward function, we examine three variants: FCESR without sparsity reward and FCESR with only factual reward. Experiments are conducted on the Baby and Beauty datasets, with results using GRU4REC reported in Table 5. Similar trends are observed across other backbone models. The findings show that each component contributes significantly to the overall performance, demonstrating the robustness of the reward design.
The results illustrate a trade-off between completeness and conciseness. On Baby, removing the feature entropy reward (w/o fe) notably degrades performance ($F_{ns}$ drops to 92.97%), while omitting the sparsity reward (w/o spars) results in a longer average length (5.64). The complete FCESR achieves the optimal balance, maintaining high accuracy ($F_{ns}$ 98.11%) with the shortest average length (5.58). A similar pattern is observed on Beauty, where the variant without sparsity reward produces the longest explanations, whereas the complete model significantly reduces length without compromising effectiveness. These results confirm the necessity of combining multiple reward components.
Although simplified variants may excel on individual metrics, the complete FCESR offers the best trade-off between explanation quality, conciseness and comprehensiveness.
@@ 5.6. Time Complexity Analysis (RQ5)
The experiments were conducted on a computational infrastructure running Ubuntu 18.04, with 8 CPU cores and 64GB of memory. Across the three datasets, the average runtime per episode is approximately 115.6 seconds on Baby, 218.2 seconds on Beauty, and 253.2 seconds on MovieLens. The total training time varies depending on the stopping criteria (reward convergence, network convergence, or maximum episode limit). The runtime performance exhibits correlations with both the average session length and the total number of sessions in each dataset, as our framework needs to traverse all items within a session during each episode. For instance, the Baby dataset exhibits notably shorter runtime compared to the other two datasets, primarily attributed to its shorter session length. Overall, the computational complexity remains moderate, with session length being the dominant factor affecting runtime performance.